OpenStudy (anonymous):

how to determine if points a(-5,-3),b(-1,-1)and c(11,5) are collinear

OpenStudy (anonymous):

collinear :three or more points that lie in a straight line

OpenStudy (anonymous):

(ab)^2 +bc^2 not equal to ca^2

OpenStudy (anonymous):

can you explain?

OpenStudy (anonymous):

Get the equation of the line using the first 2 points. Then substitute the last point in this equation to check if it is satisfied. If it is, then it means the 3 points are co-linear.

OpenStudy (anonymous):

(graphical method) can be done too, plot the points on a piece of graph paper. You should be able to tell quickly if all the points fall on a straight line. If they don't, they're noncollinear.

OpenStudy (anonymous):

how would you get the equasion of the line?

OpenStudy (anonymous):

okay an equation to the line is defined by \[y = mx+c\] m = gradient and c = intercept...do u know that?

OpenStudy (anonymous):

i know y=mx+b where m is the slope and b is the y intercept

OpenStudy (anonymous):

also the slope is 1/2

OpenStudy (anonymous):

collinear means that that the slope of two adjacent points will be equal....

OpenStudy (anonymous):

so if the slope is the same for the 3 points it is collinear?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

okay so first to get the equation to a line u need to find the "m" which is the slope/gradient so u have two points are a(-5,-3),b(-1,-1) \[m=\frac{{y_{2}-y_{1}}}{x_{2}-x_{1}}\] get the values and substitute to get the M= rather the slope

OpenStudy (anonymous):

since slope is same-they all will lie in a same straight line...

OpenStudy (anonymous):

but what if the y intcept is different? it isnt collinear then is it?

OpenStudy (anonymous):

@Sarkar what is theres another line parellel to that line.?

OpenStudy (anonymous):

if*

OpenStudy (anonymous):

so???

OpenStudy (anonymous):

isnt there a way to determine it accurately?

OpenStudy (anonymous):

they are not collinear, assume that theres an equation y=2x+3 and another one y=2x-3 they have differnt y values ( which further proves that the points dont lie on the same line)

OpenStudy (anonymous):

Check if slope(ab) = slope(ac).

OpenStudy (anonymous):

i don agree

OpenStudy (anonymous):

the slopes are equal m=1/2

OpenStudy (anonymous):

You're thining a bit wrong thush

OpenStudy (anonymous):

i agree @Ishaan94

OpenStudy (anonymous):

slope is the gradient right?

OpenStudy (anonymous):

rise over run is the slope

OpenStudy (anonymous):

Then they are collinear

OpenStudy (anonymous):

Yes, slope is gradient.

OpenStudy (anonymous):

you are confusing it@thush,you have are considering two lines,why???? the points in Q satisfy only ONE equation not 2..

OpenStudy (anonymous):

i thik the equasion is y=-x+2 where the 2 is the y intercept

OpenStudy (anonymous):

 what do u say now?